OpenStudy (anonymous):

please help The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 10 degrees Celsius is.

OpenStudy (callisto):

Solve the equation for F and show all steps. => Make F as the subject C = 5/9 (F − 32) 1. divide both sides by 5/9 \[\frac{C}{\frac{5}{9}} = \frac{\frac{5}{9} (F − 32)}{\frac{5}{9}}\]\[\frac{9C}{5} = (F − 32)\] 2. add 32 to both sides \[\frac{9C}{5} = (F − 32)\]\[\frac{9C}{5}+32 = (F − 32)+32\]\[\frac{9C}{5}+32 = F\] 3. Make F on the left \[F= \frac{9C}{5}+32\]

OpenStudy (callisto):

For part 2, put C =10 into \(F=\frac{9C}{5}+32\) \(F=\frac{9(10)}{5}+32 = ... ?\)